OpenStudy (anonymous):

A line passes throught the points (-10, -4) and (-1,2). What is the y-intercept of the line?

OpenStudy (anonymous):

1. Find the gradient of the line which passes through those two points 2. Form an equation for that line 3. Let x=0 and find the y-intercept.

OpenStudy (whpalmer4):

If you have two points \((x_1,y_1)\) and \((x_2,y_2)\), the slope \(m\) of the line passing through them is given by \[m = \frac{y_2-y_1}{x_2-x_1}\] You can then write the equation of the line using the point-slope formula: \[y-y_1 = m(x-x_1)\] The y-intercept can be found by letting \(x=0\) and solving for \(y\)